OpenStudy (anonymous):

what is the answer in the equation 4x+y=6 and y=6-4x?

OpenStudy (dumbcow):

they are the same

OpenStudy (angela210793):

wh answer? O.o

OpenStudy (anonymous):

they are the same equations

OpenStudy (anonymous):

Erm, through what process, exactly?

OpenStudy (anonymous):

subtract 4x from both sides

OpenStudy (anonymous):

so the answer is indefinite?

OpenStudy (anonymous):

yah, in a system of equation

OpenStudy (dumbcow):

indefinite? if this is a system of 2 equations then there are infinite solutions

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

I'll arrange it first. ((I followed this pattern: x + y = 6. So, in the second equation, which is y = 6 - 4x, I transposed -4x so that it will follow the pattern. Thus, negative will become positive. This is what turned out.)) 4x + y = 6 4x + y = 6 ELIMINATE X ((Since they have the both signs, the process is subtraction.)) 4x + y = 6 - 4x + y = 6 = 0x + 0y = 0 x = 0 ELIMINATE Y ((Choose just one in either of the equations. Replace the x in here with the x you came up with in the past equation. Then proceed.)) 4x + y = 6 4(0) + y = 6 0 + y = 6 y = 6 SO. x = 0 y = 6 Hope that helped.

OpenStudy (anonymous):

thank u very much alexfeb:)

OpenStudy (anonymous):

No problem. (: Though, this process is only for finding x and y through elimination. Be careful if this wasn't the thing you were looking for.

OpenStudy (dumbcow):

no thats wrong, well it is a solution but not the only solution other solutions are: x=1,y=2 x=3, y=-6 ... in elimination if you have the situation 0x+0y = 0 it means infinite solutions, the equation are the same, Not x=0